### v1

# A Novel Sampled Clustering Algorithm for Rice Phenotypic Data

Abstract
Clustering species of the same plant into different groups is an important step in developing new species of the concerned plant. Phenotypic (or Physical) characteristics of plant species are commonly used to perform clustering. In one of our recent works (Shastri et al. (2021)), we used a probabilistically sampled (using pivotal sampling) and spectrally clustered algorithm to group soybean species. These techniques were used to obtain highly accurate clusterings at a reduced cost. In this work, we extend the earlier algorithm to cluster rice species. We improve the base algorithm in three ways. First, we propose a new function to build the similarity matrix in Spectral Clustering. Commonly, a natural exponential function is used for this purpose. Based upon the spectral graph theory and the involved Cheeger’s inequality, we propose the use a base “$a$” exponential function instead. This gives a similarity matrix spectrum favorable for clustering, which we support via an eigenvalue analysis.
Second, the function used to build the similarity matrix in Spectral Clustering was earlier scaled with a fixed factor (called global scaling). Based upon the idea of Zelnik-Manor and Perona (2004), we now use a factor that varies with matrix elements (called local scaling) and works better. Third, to compute the inclusion probability of a specie in the pivotal sampling algorithm, we had earlier used the notion of deviation. This deviation for a specie cumulatively captured how far specie’s characteristic values were from their respective base values (computed over all species). A maximum function was used before to find the base values. We now use a median function, which is more intuitive. We support this choice using a statistical analysis. We address the challenge of computing the ideal number of clusters by using two complementary methods. With experiments on 1865 rice species, we demonstrate that in terms of silhouette values, our new Sampled Spectral Clustering is $61\%$ better than Hierarchical Clustering (currently prevalent). Also, our new algorithm is significantly faster than Hierarchical Clustering due to the involved sampling.
keywords:
Rice Phenotypic Data, Spectral Clustering, Spectral Graph Theory, Cheeger’s Inequality, Hierarchical Clustering, Pivotal Sampling
[…]
[label1]organization=Math of Data Science & Simulation (MODSS) Lab, Computer Science & Engineering,addressline=IIT Indore, city=Indore, postcode=453552, state=Madhya Pradesh, country=India
[…]
[label3]organization=ICAR-Indian Institute of Soybean Research,city=Indore, postcode=4520001, state=Madhya Pradesh, country=India
1 Introduction
Phenotypic characteristics (or physical characteristics) of plant species are often used in clustering them into separate categories (Painkra et al. (2018), Sharma et al. (2014)). This is done so that plant species from different categories (or diverse plant species) could be selectively chosen for developing new species having better characteristics (Swarup et al. (2021)) (or called breeding). Hierarchical Clustering is one of the most commonly used clustering algorithms in this domain (Kahraman et al. (2014)). This algorithm suffers from low accuracy issues and is also computationally expensive. In one of our recent works, we used Spectral Clustering instead (Shastri et al. (2021)), which is considered to be one of the most accurate clustering algorithms. Since Spectral Clustering, like Hierarchical Clustering, is computationally costly, we sampled the plant species before clustering. Pivotal sampling, which is probability based, was used due to its high accuracy and low cost (requires a single pass through the data). Pivotal sampling has the added advantage of maintaining independence from the data’s order or location.
In our earlier work (as discussed above), we applied sampled clustering to soybean plant species. In this current work, we focus on rice plant species. Although our earlier algorithm for soybean can be directly applied to rice, we improve it further in this work as well. We have three novel contributions.
In the spectral clustering algorithm, one of the crucial step is building the similarity matrix. If $d_{p_{1}p_{2}}$ denotes the distance between two plant species $p_{1}$ and $p_{2}$, then the element $W_{p_{1}p_{2}}$ of the similarity matrix $W$ should be inversely proportional to $d_{p_{1}p_{2}}$ (because larger the distance between two species, smaller should be the similarity between them).
In our earlier work, we had taken $W_{p_{1}p_{2}}$ to be the most commonly used choice of $e^{-\frac{d_{p_{1}p_{2}}}{2\sigma^{2}}}$, where $\sigma$ defines the decay of the distance. In this work, we propose a base “$a$” exponential function instead of a natural exponential. This results in similarity matrix spectrum supportive of better clustering (theorized by the spectral graph theory, specifically the Cheeger’s inequality). We use $W_{p_{1}p_{2}}$ as $30^{-\frac{d_{p_{1}p_{2}}}{2\sigma^{2}}}$. We justify this choice with an eigenvalue analysis. This is our first novelty.
Having a global decay factor ($\sigma$) as above is not ideal. In our implementation, we use the concept of local scaling (Zelnik-Manor and Perona (2004)), which according to the theory works better. That is, $W_{p_{1}p_{2}}$ is taken as $30^{-\frac{d_{p_{1}p_{2}}}{\sigma_{p_{1}}\sigma_{p_{2}}}}$, where $\sigma_{p_{1}}$ and $\sigma_{p_{2}}$ are computed via a K-nearest neighbour (KNN) search. This is our second novelty.
Computing probabilities in the pivotal sampling algorithm is an important step. To compute these probabilities, we introduce the notion of deviation. If $(\kappa_{j})_{i}$ represents the value of the $j^{th}$ characteristic for the $i^{th}$ specie, we define total deviation for the $i^{th}$ specie as
$dev_{i}=\sum_{j=1}^{m}(base((\kappa_{j}){}_{i=1,...,n})-(\kappa_{j}){}_{i}),$ (1)
where $m$ denotes the total number of the characteristics and the base function captures the representative value of the $j^{th}$ characteristic over all the species ($n$). A specie with a lower deviation is more likely to represent the entire population. Thus, we define the inclusion probability of a specie to be inversely proportional to this deviation. In our earlier work, we had taken base to be the maximum function. In this work, we propose the use of a median function instead, which is more intuitive. We support this choice via a statistical analysis. This is our third novelty.
[…]
The rest of the paper has six more sections. In Section 2, we discuss the past work in this domain. Section 3 presents both our old and new Spectral Clustering. In Section 4, we discuss pivotal sampling and its application. Section 5 describes the setup for data and computation of the ideal number of clusters. In Section 6, we analyze the design choices of our new sampled Spectral Clustering. Section 7 presents the results. Finally, in Section 8, we provide conclusions and future work.
2 Literature Survey
Here, we discuss relevant past work done in clustering of rice species using phenotypic data. We do not plan to be exhaustive.
Islam et al. (2020) employed the Unweighted Pair Group Method with Arithmetic Mean (UPGMA), which is a type of Hierarchical Clustering. The paper investigated clustering patterns of 10 species of Thai indigenous upland rice using phenotypic traits. Cluster analysis resulted in three distinct groups. This highlighted the potential for utilizing traits such as flag leaf length, harvest index, total dry weight, total grain weight, and filled grains in obtaining higher grain yield.
Verma et al. (2019) used UPGMA-based dendrogram analysis, also a type of Hierarchical Clustering. The paper investigated the clustering of 114 rice species of North East India using phenotypic traits. The species were clustered into seven groups. Grouping of rice species was done on $15$ traits which include root length, shoot length, fresh root length, dry root length, root volume, root angle, peripheri root, etc. This was used for the assessment of genetic variation in rice breeding programmes.
Islam et al. (2018) employed Neighbor-joining tree analysis, which is a type of Hierarchical Clustering. In this paper, the authors clustered $15$ red rice varieties in Bangladesh using phenotypic traits. The species were clustered into five groups (I, II, III, IV, V). A total of 15 traits, e.g., culm diameter, plant height, panicle length, five panicle weight, filled grains per panicle, etc. were used for the detection of economically desired traits for future molecular breeding programmes.
Pachauri et al. (2013) used UPGMA for Clustering. The paper investigated the clustering of 41 rice species cultivated by Indian farmers using phenotypic traits. The species were clustered into four groups. The phenotypic traits, such as the number of grains per panicle, thousand-grain weight, and the measurements of length and breadth for both whole and milled grains, were used in breeding programs.
All the above works have two major drawbacks. First, they used Hierarchical Clustering, which is not very accurate and computationally expensive. Second, they experimented on small datasets (a few tens to a few hundreds of species).
We overcome both these drawbacks here. As mentioned in the introduction, we use Spectral Clustering (old from Shastri et al. (2021) and new), which is known to be highly accurate. To offset the high computational cost we use sampling (old from Shastri et al. (2021) and new). Also, we experiment with about $2,000$ species of rice.
3 Spectral Clustering and A New Variant
In this section, we discuss three things. First, we present the standard spectral clustering algorithm. Next, we drive our novel locally-scaled base “$a$” exponential spectral clustering algorithm. Finally, we discuss the application of both to rice phenotypic data.
3.1 Standard Spectral Clustering
Spectral Clustering is one of the most popular modern clustering algorithms. It is simple to implement and can be solved efficiently by standard linear algebra software. Given a set of points $S=\{p_{1},p_{2},...,p_{n}\}$ in $R^{m}$ that we want to cluster into k subsets, the algorithm consists of below steps Von Luxburg (2007).
  1. 1.
    Form a similarity matrix A such that
    $A_{i}{}_{j}=e^{\left(-\frac{d_{p_{i}p_{j}}}{2\sigma^{2}}\right)},$ (2)
    with $i$, $j$in $\{1,...,n\}$ and $A_{ii}=0$. Here, $d_{p_{i}p_{j}}$ denotes the distance between two points $p_{i}$ and $p_{j}$ and $\sigma$ defines the decay of the distance.
  2. 2.
    Construct the Laplacian matrix
    $L=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}},$ (3)
    where $D$ is a diagonal matrix whose $(i,i)$ element is the sum of the elements of $A$’s $i^{t}{}^{h}$ row.
  3. 3.
    Let $e_{1}$, $e_{2}$., $e_{k}$ be the first $k$ eigenvectors of $L$. Then, form the matrix $X=[e_{1},e_{2}....,e_{k}]$ by stacking the eigenvectors as columns of this matrix.
  4. 4.
    Form $Y$ by normalizing $X$’s rows to unit length.
  5. 5.
    Cluster $Y$ using the $k-$Means clustering.
The earlier work that we extend here (Shastri et al. (2021)) used this standard Spectral Clustering.
3.2 Locally Scaled Base “$a$” Exponential Function Spectral Clustering
Spectral Clustering is based upon spectral graph theory. To derive our new algorithm, we first revisit few concepts from this domain. We form a graph from the given data as follows (Gharan (2020)): (a) use data points as vertices and, (b) connect each point with the remaining points with an edge having weight equal to the corresponding element of similarity matrix A.
Definition 3.1 (Conductance (Gharan (2015))).
Given a graph $G=(V,E)$ with $V$ partitioned into $S$ and $\overline{S}$, the conductance of $S$ is defined as
$\phi(S)=\frac{|E(S,\overline{S})|}{Vol(S)},$ (4)
where numerator is the fraction of edges in $cut(S,\overline{S})$ and denominator is the sum of vertices in $S$. The conductance of $G$ is defined as
[…]
or the smallest conductance among all sets with at most half of the total volume.
Theorem 3.1 (Cheeger’s Inequality (Gharan (2015))).
For any graph $G$,
$\frac{\lambda_{2}}{2}\leq\phi(G)\leq\sqrt{2\lambda_{2}},$ (6)
where $\lambda_{2}$ is the $2^{nd}$ smallest eigenvalue of $L$ given by (3).
From the above theorem, we infer that $\phi(G)$ is close to zero (or G can be grouped into $2$ clusters) if and only if $\lambda_{2}$ is close to zero. Note that $\lambda_{1}$ is always zero. This characterization carries over to higher multiplicities as well. $G$ can be grouped into $k$ clusters if and only if there are $k$ eigenvalues close to zero (Lee et al. (2014)).
We propose using a base “$a$” exponential function instead of the natural exponential function in (2) of the standard spectral clustering algorithm. That is,
$A_{ij}=a^{\left(-\frac{d_{p_{i}p_{j}}}{2\sigma^{2}}\right)},$ (7)
where “$a$$>$$e$”. This results in $A_{ij}$ of (7) being smaller than $A_{ij}$ of (2).
Theorem 3.2.
The elements of non-normalized Laplacian matrix $L=D-A$ get smaller in absolute sense when we use (7) instead of (2), with “$a$$>$$e$”, to build A. Here, $D$ is the diagonal matrix whose $(i,i)$ element is the sum of $i^{th}$ row of $A$. Further, this leads to reduction in upper bound of eigenvalues of L.
Proof.
The first part of the Theorem is obvious. Since elements of $A$ get smaller with the proposed change of base, the elements of $D$ also get smaller ($D$ is formed via elements of A). Thus, elements of $D-A$ or $L$ get smaller in the absolute sense. For the second part of the proof, we use the fact that the spectral radius of the matrix is bounded above by its norm or $\rho(L)\leq||L||$. ∎
Conjecture 3.3.
The above theorem holds true when we change the non-normalized Laplacian matrix $L=D-A$ with the normalized Laplacian matrix $L=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$.
We are unable to prove this theoretically. However, this holds true experimentally. We demonstrate in the Analysis section later in this paper that the change of the base as discussed in the above conjecture leads to a reduction in the eigenvalues of $L$. Thus, from the Cheegers’s Inequality (6), we infer that we get a better clustering when we use base “$a$” exponential function instead of the natural exponential function in building the similarity matrix. This is supported by experiments in the Results section.
Next, to further improve our clustering, we depart from the conventional practice of utilizing a global scaling factor ($\sigma$) in (7). Instead, we adopt the concept of a local scaling factor specific to each data point, as proposed by Zelnik-Manor and Perona (2004). Now, the similarity between the two points is defined as
$A_{ij}=a^{\left(-\frac{d_{p_{i}p_{j}}}{\sigma_{i}\sigma_{j}}\right)}.$ (8)
The determination of the local scale $\sigma_{i}$ involves analyzing the local statistics within the neighborhood of a given point. We employ a simple yet effective approach for scale selection. That is,
$\sigma_{i}=d_{p_{i}p_{K}},$ (9)
where $p_{K}$ is the $K^{\text{th}}$ neighbor of $p_{i}$. The selection of K is independent of the scale and based upon the data dimensionality. Our experiments demonstrate the superiority of local scaling.
3.3 Application to Rice Phenotypic Data
Let us consider a dataset consisting of $n$ species with $m$ distinct characteristics. We begin by normalizing the characteristics as follows (Shastri et al. (2021)):
$(\chi_{j}){}_{i}=\frac{(x_{j}){}_{i}-\min(x_{j})}{\max(x_{j})-\min(x_{j})}.$ (10)
[…]
$p_{i}=\begin{bmatrix}(\chi_{1}){}_{i}\\ (\chi_{2}){}_{i}\\ \vdots\\ (\chi_{m}){}_{i}\\ \end{bmatrix},$
for $i=1,2,..,n$.
We compute the distance between specie $p_{i}$ and $p_{j}$ in (8), i.e., $d_{p_{i}p_{j}}$ using Euclidean, Squared-Euclidean, and Correlation distance measures. These measures work out best for us and are given below.
  • 1.
    Euclidean: It represents the straight-line distance between two points in Euclidean space, and is calculated as follows:
    $d_{ij}=\sqrt{\sum_{l=1}^{m}(p_{i}^{l}-p_{j}^{l})^{2}},$ (11)
    where $p_{i}^{l}$ is $(\chi_{l}){}_{i}$ and $p_{j}^{l}$ is $(\chi_{l}){}_{j}$.
  • 2.
    Squared-Euclidean: It is the square of the Euclidean distance, and is given as follows:
    $d_{ij}=\sum_{l=1}^{m}(p_{i}^{l}-p_{j}^{l})^{2},$ (12)
    with $p_{i}^{l}$ and $p_{j}^{l}$ are defined as above.
  • 3.
    Correlation: It captures the correlation between two non-zero vectors, and is expressed as follows:
    $d_{ij}=1-\frac{(p_{i}-\bar{p_{i}})^{t}(p_{j}-\bar{p_{j}})}{\sqrt{(p_{i}-\bar{p_{i}})^{t}(p_{i}-\bar{p_{i}})}\sqrt{(p_{j}-\bar{p_{j}})^{t}(p_{j}-\bar{p_{j}})}},$ (13)
    where $\bar{p_{i}}$ and $\bar{p_{j}}$ represent the means of vectors $p_{i}$ and $p_{j}$, respectively, multiplied by a vector of ones, and the $t$ indicates the transpose operation.
4 Pivotal Sampling and its Application
Pivotal Sampling (Deville and Tille (1998)) is an effective non-uniform probabilistic sampling technique without replacement, which employs a fixed sample size. This technique involves sequentially scanning of the population, wherein at each iteration there is a duel between two units, $i$ and $j$, with inclusion probabilities of $\pi_{i}$ and $\pi_{j}$, respectively, to get selected or rejected from the sample.
  • 1.
    Selection step: If $(\pi_{i}+\pi_{j}\geq 1)$, then one of the units will be selected (i.e., sampled).
  • 1.
    Rejection step: If $(\pi_{i}+\pi_{j}<1)$, then one of the units will be rejected (i.e., not sampled).
In our earlier work, the probabilities in the pivotal sampling algorithm were taken as a function of variations in the characteristics of the plant species. Plant height, number of pods per plant, days to pod initiation, seed yield per plant, etc., are some characteristics examples. Specifically, this probability for the $i^{th}$ specie was taken to be inversely proportional to
$dev_{i}=\sum_{j=1}^{m}((max(\chi_{j}){}_{i=1,..,n})-(\chi_{j}){}_{i}),$ (14)
where $dev_{i}$ denotes the deviation for the $i^{th}$ specie, $max(\chi_{j})$ denotes the maximum value of the $j^{th}$ characteristic over all the species, and $(\chi_{j})_{i}$ denotes the value of $j^{th}$ characteristic for the $i^{th}$ specie.
This method of computing deviation implies that the higher the characteristic values of a specie, the higher its probability. Thus, the higher chance that it would be selected in the sample (and would represent all the species). However, species with higher characteristic values representing all species is not ideal. Take ‘days to pod initiation’ characteristic for example. Here, it could be that most species take less number of days to pod initiation. Hence, having a specie that takes more number of days for pod initiation, to represent all would be inaccurate.
Hence, we propose the use of the median function instead of the maximum function in (14). This would ensure that the samples consist of species with characteristic values closer to the median values. Thus, such sampled species would represent all the species better. We support our this choice by performing statistical analysis as well.
The final inclusion probability of $i^{th}$ specie is obtained by normalizing as follows (Shastri et al. (2021)):
$\pi_{i}=N\frac{\frac{1}{dev_{i}}}{\sum_{i=1}^{n}\frac{1}{dev_{i}}}.$ (15)
We then iteratively apply the selection and rejection steps of the pivotal sampling technique in order to obtain a sample of size $N$.
Next, we perform earlier discussed clustering techniques on the sampled data of size $N$. Since the total data is of size $n$ (with $n>N$), there is a need to reverse map the remaining $n-N$ data points to the required cluster. For this, we define the notion of average similarity, which between the non-clustered specie $\tilde{p}$ and the cluster $C_{k}$ is given as
$AS(C_{k},\tilde{p})=\frac{1}{\#(C_{k})}\sum_{y\in C_{k}}a^{\left(-\frac{{}^{d}\tilde{p}y}{\sigma_{\tilde{p}}\sigma_{y}}\right)}.$ (16)
Here, #$\#(C_{k})$ represents the number of species in cluster $C_{k}$. We compute the average similarity of $\tilde{p}$ with all the $k$ clusters and assign it to the cluster with which $\tilde{p}$ has the maximum similarity.
5 Setup
Few settings of our algorithm from Section 3.2 are as follows: (a) The best value of “$a$” (the base of the exponential function used to build the similarity matrix) for us turns to be “$30$”. (b) The most fitting value of $K$ (neighbor of a point in local scaling) for the case of clustering without sampling comes to $180$, and for the case of sampled clustering it comes to be $400$. In rest of this section, we initially describe the rice data, and then we discuss different ways of computing the ideal number of clusters.
5.1 Data Description
Our algorithm can be applied to any plant data, however, in this work, we do experiments on the rice data. This data is taken from The International Rice Information System (IRIS) (www.iris.irri.org)- a platform for meta-analysis of rice crop plant data (McLaren et al. (2005)). It consists of 12 phenotypic (or physical) characteristics of $1865$ rice species. A snapshot of this data is given in the Table 1.
Sr. No. Cudicle Cultural Cuneiform Grain Grain Grain
Repro Repro Repro Length Width Weight
-duction -duction -duction per
100 Seed
1 5 147 16 8.7 3.1 2.9
2 6 150 27 7.1 3.3 2.1
1875 3 56 16 7.7 3.4 2.8
Sr. No. HDG Lightness Leaf Leaf Plant Post Stem
80HEAD of Color Length Width Harvest Height
Traits
1 102 25 72 1.1 29 54
2 123 20 73 1.5 27 45
1875 69 10 31 1 16 23
Table 1: Phenotypic data of rice plant.
5.2 Ideal Number of Clusters
Figure 1: First 30 eigenvalues obtained using natural exponential function and base “$30$” exponential function for building the similarity matrix.
Determining the ideal number of clusters is always a dataset specific task, and we have many methods to approximate this number. A commonly used method is the Eigen-Gap heuristic (Von Luxburg (2007)) tailored for Spectral Clustering. This heuristic says that given $\lambda_{i}$’s as the eigenvalues of the Laplacian matrix, $k$ is the ideal number of clusters when the eigenvalues $\lambda_{1}$, $\lambda_{2}$, .., $\lambda_{k}$ have small magnitude while $\lambda_{k+1}$ has comparatively much higher magnitude. Figure 1 displays the $30$ smallest eigenvalues derived from the Laplacian matrix obtained during the application of our locally scaled base “$30$” exponential Spectral Clustering on the rice data. From this figure, we see a jump at 1, 2, 3, 5, and 10. To double-check our result, we also apply the Spectral Global Silhouette method proposed in Mur et al. (2016). This method uses the concept of local scaling, as discussed earlier, and gives 5 as the ideal number of clusters. Plant biologists are typically interested in more number of clusters, and hence, we choose 5, 10, 15, and 20 as the ideal number of clusters for our experimentation.
6 Analysis
In this section, first, we demonstrate that eigenvalues of the Laplacian in (3) reduce when we use an exponential function with base $a$ (i.e., (7)), instead of a natural exponential function (i.e., (2)) in building the similarity matrix. Second, we check the quality of our samples via a sampling analysis.
6.1 Eigenvalue Analysis
Table 2 lists the first 30 smallest eigenvalues of the Laplacian matrix obtained from both the natural exponential function similarity matrix and base “$30$” exponential function similarity matrix. Additionally, Figure 1 plots these eigenvalues. From both the table and the figure, we observe that the eigenvalues associated with the latter similarity matrix are closer to zero as compared to the eigenvalues associated with the former similarity matrix. This finding validates our Conjecture 3.3. Thus, using Cheeger’s Inequality (8), we infer that our novel approach of constructing the similarity matrix using a base “$30$” exponential function should lead to improved clustering. This is supported experimentally in the next section.
Sr. No. 1 2 3 4 5 6 7 8 9
Base (e) Fun 0.0000 0.6653 0.8397 0.8677 0.9038 0.9159 0.9235 0.9402 0.9472
Base (30) Fun 0.0000 0.4272 0.6642 0.7078 0.7134 0.8025 0.8151 0.8261 0.8441
Sr. No. 10 11 25 26 27 28 29 30
Base (e) Fun 0.9538 0.9606 0.9916 0.9918 0.9926 0.9929 0.9934 0.9936
Base (30) Fun 0.8468 0.8761 0.9420 0.9439 0.9457 0.9468 0.9501 0.9506
Table 2: Small eigenvalues of Laplacian matrices obtained using different similarity matrices.
6.2 Sampling Analysis
In order to assess the quality of the obtained samples, we employ the estimation of the following metrics: the Population Total (PT) estimator and the Horvitz-Thompson (HT) estimator. We define the PT estimator for a specific characteristic as the summation of all the values associated with that characteristic. That is, for a characteristic $j$, it is computed as
$(Y_{PT})_{j}=\sum_{i=1}^{U}(\chi_{j}){}_{i},$ (17)
where U is the set of all species (or #(U)=n) and $(\chi_{j}){}_{i}$ is the value of the $j^{t}{}^{h}$ characteristic for the $i^{t}{}^{h}$ specie.
The HT estimator is defined as follows:
$(Y_{HT})_{j}=\sum_{i=1}^{S}\frac{(\chi_{j}){}_{i}}{\pi_{i}},$ (18)
where $S$ is the set of samples (or #(S)=N), $(\chi_{j}){}_{i}$ is defined as above, and $\pi_{i}$ denotes inclusion probability of $i^{t}{}^{h}$ specie. The sampling is better when the values of the HT estimators are closer to the PT estimators.
Characteristics PT HT (Max) HT (Median)
estimators estimators estimators
  1 9.037800e+03 9.099514e+03 9.032072e+03
  2 2.121750e+05 2.113126e+05 2.103986e+05
  3 3.057500e+04 3.120878e+04 3.089743e+04
  4 1.614470e+04 1.619068e+04 1.612586e+04
  5 5.661400e+03 5.625295e+03 5.652058e+03
  6 4.630100e+03 4.599130e+03 4.538882e+03
  7 1.865000e+05 1.885044e+05 1.852332e+05
  8 3.343930e+04 3.508258e+04 3.456751e+04
  9 1.024092e+05 1.034117e+05 1.026595e+05
  10 2.569800e+03 2.552394e+03 2.562038e+03
  11 4.660790e+04 4.639347e+04 4.629499e+04
  12 7.214200e+04 7.184603e+04 7.227474e+04
Table 3: PT and HT Estimators.
Table 3 lists the PT estimators, HT estimators when using the maximum function in (14), and HT estimators when using the median function in (14) for all the 12 characteristics. The first thing we observe from the table is that the PT estimators are close to both the sets of HT estimators, and hence, both the samplings are good. Second, we also observe that the PT estimators are closer to the HT estimators when using the median function than the HT estimators when using the maximum function.
To better demonstrate the second observation, we plot the difference between the PT and the HT estimators (when using both the maximum function and the median function) for all characteristics in Figure 2. We can notice from this figure that the median function leads to a smaller difference in 9 of the 12 characteristics as compared to the maximum function. Thus, median based sampling is better.
Figure 2: Difference between PT and HT estimator when sampled with maximum function and median function.
Old New
Clusters Distance Old SC New SC HC Vs Vs
HC % HC %
Euclidean 0.1249 0.1258 0.07
5 SqEuclidean 0.215 0.2242 0.1349 28.88 34.15
Correlation 0.22 0.229 0.1707
Euclidean 0.0952 0.1113 0.0662
10 SqEuclidean 0.1592 0.2002 0.1079 35.64 55.43
Correlation 0.1747 0.1987 0.1288
Euclidean 0.0841 0.0946 0.026
15 SqEuclidean 0.1342 0.1725 0.0677 51.55 72.33
Correlation 0.1517 0.1719 0.1001
Euclidean 0.0784 0.0874 0.0128
20 SqEuclidean 0.1109 0.1587 0.0432 83.35 103.40
Correlation 0.1454 0.1613 0.0793
Average percentage gain 49.86 66.33
Table 4: Silhouette values of clustering algorithms.
7 Results
In this section, we perform two kinds of experiments. In Section 7.1, we compare the different clustering algorithms without sampling. In Section 7.2, we compare the sampled clustering algorithms.
7.1 Clustering without Sampling
Here, we compare three clusterings. First is the standard Spectral Clustering as described in Section 3.1 (also natural exponential Spectral Clustering), and used in Shastri et al. (2021). We refer to this as Old SC. Second is the locally scaled base “$30$” exponential Spectral Clustering as elaborated in Section 3.2. We call this the New SC. Finally, the third is Hierarchical Clustering, which is mentioned in Section 2. We refer to this as HC.
The results of this comparison are given in Table 4. Here, the first column denotes the number of the clusters that are chosen based upon the previous analysis. The second column contains the distance metrics used to build the similarity matrix in the spectral clustering algorithms. Columns three through five list the silhouette values of the respective algorithms. Best values in a cell are highlighted in bold. Finally, columns six and seven give the percentage gain of Old SC and New SC over HC, respectively. The best values are used to compute this gain. As evident, the New SC works the best here.
Old New Sampled Old New
Clusters Distance Sampled SC Sampled SC HC Vs Vs
HC % HC %
Euclidean 0.1218 0.127 0.1086
5 SqEuclidean 0.2111 0.2233 0.1426 25.61 30.08
Correlation 0.2276 0.2357 0.1812
Euclidean 0.0851 0.1043 0.0664
10 SqEuclidean 0.1602 0.1943 0.1 26.20 45.68
Correlation 0.171 0.1974 0.1355
Euclidean 0.0815 0.0858 0.0591
15 SqEuclidean 0.1348 0.1586 0.0917 11.01 15.10
Correlation 0.1573 0.1631 0.1417
Euclidean 0.0751 0.0774 0.0453
20 SqEuclidean 0.1115 0.1367 0.0856 6.68 14.81
Correlation 0.1405 0.1512 0.1317
Average percentage gain 17.37 26.42
Table 5: Silhouette values of clustering algorithms with sampling.
Clusters Distance New HC New Sampled
Sampled SC SC vs HC %
Euclidean 0.127 0.07
5 SqEuclidean 0.2233 0.1349 38.08
Correlation 0.2357 0.1707
Euclidean 0.1043 0.0662
10 SqEuclidean 0.1943 0.1079 53.26
Correlation 0.1974 0.1288
Euclidean 0.0858 0.026
15 SqEuclidean 0.1586 0.0677 62.93
Correlation 0.1631 0.1001
Euclidean 0.0774 0.0128
20 SqEuclidean 0.1367 0.0432 90.67
Correlation 0.1512 0.0793
Average percentage gain 61.24
Table 6: Silhouette values of New Sampled SC and HC.
7.2 Sampled Clustering
Here, we again compare the three clusterings (as in the previous section) but now after sampling, First, is the standard Spectral Clustering from Section 3.1 (natural exponential Spectral Clustering) along with the maximum function based Pivotal Sampling from Section 4, as used in Shastri et al. (2021). We refer to this as Old Sampled SC. Second is the locally scaled base “$30$” exponential Spectral Clustering from Section 3.2 along with the median function based Pivotal Sampling from Section 4. We call this New Sampled SC. Finally, as mentioned earlier, the third is the Hierarchical Clustering along with the median based Pivotal Sampling. We refer to this as Sampled HC. The results for this are given in Table 5. All columns here are the same as in Table 4. As evident, the New Sampled SC works best again.
To conclude, in Table 6, we compare New Sampled SC with HC. Thus, our new algorithm is substantially better than the current best and also computational cheaper.
8 Conclusions and Future Work
In one of our recent works, we had developed a probabilistically sampled (using pivotal sampling) and spectrally clustered algorithm for grouping soybean species. In this work, we improve this algorithm and apply it to rice species. The first two improvements involve use of better functions to build the similarity matrix in the spectral clustering algorithm. These design choices are supported via an eigenvalue analysis. The third improvement is using a more intuitive function (median) instead of a maximum function in computing the probabilities in the pivotal sampling algorithm. This design choice is supported via the computation of statistical measures.
[…]
There are multiple future work directions here. First, in one of the seminal works (Ng et al. (2001)), the authors have listed sufficiency conditions for SC to work well. It would be very useful to translate those conditions to plant data. Second, it would be useful to experiment with other accurate clusterings and samplings (e.g., see Jain et al. (2021)). Third, although phenotypic characteristics are useful for clustering, genetic data of plant species carries more information. In our earlier work (Shastri et al. (2019)), we had explored the possibility of using genetic data for clustering, however, reduced data was used there. It would be interesting to experiment with the full data exhaustively.